Suppose that you have two possible projects. Project A gives a profit of 10 in period 0 and 9 in period 1. Project B gives a profit of 12 in period 0 and 6.5 in period 1. Suppose that the interest rate r = 0.05. Which project is better? (Hint: no matter when you want to consume, higher PDV is always better.)
Answer #1

Present discounted value of Project A = 10 + 9/(1+ 0.05)

= 10 + 8.57 = 18.57

Present discounted value of Project A = 12 + 6.5/(1+ 0.05)

= 12 + 6.19 = 18.19

Since PDV of project A is higher, it is the better project